OpenStudy (anonymous):

find a if (a,4) and (2,5) are parallel vectors ?

OpenStudy (savvy):

a = 8/5

OpenStudy (anonymous):

can u show the working ??? @Savvy

OpenStudy (savvy):

sure, since they are parallel vectors(NOTE: they are to be specified as position vectors) hence the co-ordinates of the vectors should be proportional.... 5/4 = 2/a

OpenStudy (savvy):

or else you could find a unit vector in the direction of the 2nd vector and multiply it by the required magnitude which can be deciphered by the y co-ordinate of the 1st vector...